OpenStudy (anonymous):

Given that f(x) = 4x − 3 and g(x) = 2x - 1 / 3 solve for g(f(2)). 1 3 9 15

OpenStudy (solomonzelman):

when you have g(f)( 2)... for example: \(\normalsize\color{black}{f((g)3) }\) \(\normalsize\color{black}{f(x)=2x+5 }\) and \(\normalsize\color{black}{f(g)=4x-9 }\) \(\normalsize\color{black}{f(g)=2\color{red}{(4x-9)} +5 }\) \(\normalsize\color{black}{f(g)=8x-18+5 }\) \(\normalsize\color{black}{f(g)=8x-13 }\) \(\normalsize\color{black}{f((g)2)=8\color{red}{(3)}-13 }\) \(\normalsize\color{black}{f((g)2)=24-13 }\) \(\normalsize\color{black}{f((g)2)=11 }\)

OpenStudy (anonymous):

So what do I do to find the answer of this equation?

OpenStudy (solomonzelman):

\(\normalsize\color{black}{f(x) = 4x − 3 }\) 1) you plug in what f(g) is equal to for x. In other words, you plug in ` 2 x - 1 / 3 ` for x. 2) then after simplifying it, you plug in 2 for x, and calculate it away to find the answer.

OpenStudy (anonymous):

Okay I think I got it, can you walk me through?

OpenStudy (anonymous):

I get stuck on equations like this often:(

OpenStudy (solomonzelman):

take a shot.. if you have an error, I'll correct you .

OpenStudy (anonymous):

Okay!

OpenStudy (anonymous):

So f(2x-1/3) is 1/3f(2x-1)

OpenStudy (anonymous):

Like that?

OpenStudy (anonymous):

I simplified the first part

OpenStudy (solomonzelman):

hold on, in your f(g) is it 2x-1 ---- 3 right ?

OpenStudy (anonymous):

Yes

OpenStudy (solomonzelman):

Okay, was making sure.

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \bf f(x) = 4x − 3 }\) \(\Huge\color{blue}{ \bf g(x)=\frac{2x-1}{3} }\) \(\Huge\color{blue}{ \bf f(g)=4(\frac{2x-1}{3})-3 }\)

OpenStudy (solomonzelman):

like this. Sorry if the letters are too big

OpenStudy (anonymous):

Okk!

OpenStudy (anonymous):

Now i plug in two?

OpenStudy (solomonzelman):

first you simplify it \(\Huge\color{red}{ \rm f(g)=4(\frac{2x-1}{3})-3 }\) \(\Huge\color{red}{ \rm f(g)=\frac{8x-4}{3}-3 }\)

OpenStudy (solomonzelman):

now, you plug in 2 for x.

OpenStudy (anonymous):

Ok, so 8(2)-4/3 which is 16-4/3, which is 12/3 = 4

OpenStudy (anonymous):

4 - 3 = 1

OpenStudy (anonymous):

Is it one?

OpenStudy (solomonzelman):

sorry I got disconnected. (8 × 2 - 4) ÷ 3 - 3 (16 - 4) ÷ 3 - 3 (12) ÷ 3 - 3 4 - 3 \(\normalsize\color{black}{\underline{ 1 }}\)

OpenStudy (anonymous):

YAYYY

OpenStudy (anonymous):

Thanks so much for all the help, and teaching me!

OpenStudy (solomonzelman):

Anytime:) That's OS is for; enjoy !